OpenStudy (dtan5457):

quick trig check find values of sin2a, cos2a, and tan2a when sina=-2/3 and this is in the 3rd quadrant

OpenStudy (fibonaccichick666):

2 ways there, use double angle formulas or solve for a then plug and chug

OpenStudy (dtan5457):

I'm getting \[\sin 2a=\frac{ 4\sqrt{5} }{ 9 }\] cos2a=\[\frac{ 1 }{ 9 }\] tan2a=\[4\sqrt{5}\]

OpenStudy (fibonaccichick666):

so yea third side is (\sqrt 5\) I agree with sine

OpenStudy (fibonaccichick666):

h/o for the other two

OpenStudy (fibonaccichick666):

I agree with cosine

OpenStudy (dtan5457):

and tangent?

OpenStudy (fibonaccichick666):

and I agree for tangent as well

OpenStudy (fibonaccichick666):

looks good to me, make sure you define the third side and show how you got it in your work

OpenStudy (dtan5457):

yup, did that. thanks for checking.

OpenStudy (fibonaccichick666):

np, good job